# FaithLM: Towards Faithful Explanations for Large Language Models

### v3

@@ 4.5 Ablation Studies on Contrary Explanation (RQ3)
The quality of contrary explanations $\neg\mathcal{E}_{\text{NL}}$ determines the efficacy of FaithLM. We leverage the powerful LLMs as the LLM agent to generate contrary explanations, requesting the delivery of high-quality opposite-meaning outputs from their original NL explanations. In this section, we evaluate the quality of contrary explanations, aiming to observe the semantic differences between the original NL explanations and their contrary explanations. To examine the quality, we employ the GPT classifier and two well-trained NLI classifiers, Roberta-Large and XLNet-Large [31]. We leverage each classifier to distinguish whether the relationship between the “original NL explanations" and “contrary explanations" belong to the category of “similar meaning (entailment)," “dissimilar meaning (contradiction)," or “non-relevant (neutral)." We follow the evaluation settings from [23] on GPT-classifier with evaluation prompt provided in Table 7 of Appendix F.
@@
The results are shown in Figure 6 under the randomly sampled 100 instances from the ECQA and COPA datasets. We observe that the two NLI classifiers achieve up to 86% and 82% in the “dissimilar meanings"" category on the ECQA and COPA datasets, respectively. The empirical results of the GPT-classifier demonstrate that the derived explanations from SelfExp obtain more non-faithful information than FaithLM, which may increase the risk for the LLM agent of Fidelity Evaluator in generating non-relevant information as the contrary explanations. In addition, we provide case studies in Appendix I to demonstrate the informativeness and readability of contrary explanations.
@@ 5 Conclusion
In this paper, we introduce FaithLM to explain the decision-making process of LLMs. Specifically, FaithLM employs a fidelity enhancement strategy to progressively refine the fidelity of derived explanations and explanation trigger prompts. Experimental results on three datasets demonstrate the effectiveness of FaithLM in terms of the high fidelity of derived explanations, and better alignment with the ground-truth explanations of the datasets. As for future work, we plan to expand FaithLM to investigate high-stack applications, such as the healthcare domain, where transparency concerns are increasingly prominent due to the widespread adoption of black-box LLMs.
@@
Appendix A Details about Datasets
@@
  • ECQA [2]. ECQA is an extension of the CQA dataset [37]. Specifically, based on the CQA dataset, it annotates the positive or negative properties and golden explanations for the QA pairs. Due to API cost budgets, we evaluate our framework on the first 500 instances in the ECQA dataset.
@@
  • TriviaQA LongBench [19]. TriviaQA LongBench (TriviaQA-Long) is a reading comprehension dataset. It includes 300 question-answer-evidence triples sourced from the Longbench [5] dataset [^1] https://huggingface.co/datasets/THUDM/LongBench/. This dataset features question-answer pairs crafted by trivia enthusiasts, accompanied by independently sourced evidence documents, providing supervision for answering these questions.
@@
  • Balanced COPA [35, 20]. The Balanced COPA (COPA) dataset is a collection of 500 questions for commonsense causal reasoning. Each question consists of a premise and two alternatives, where the task is to select the alternative that more plausibly has a causal relation with the premise.
Appendix B Related Work
B.1 Post-hoc Explanation
@@
Post-hoc explanation techniques have undergone significant development and discussion, driven by the widespread adoption of black-box ML models across various data modalities. A multitude of post-hoc algorithms has been introduced from two aspects: local and global explanations [30, 13]. Explanations aim to explain the reasoning behind an individual model for each input instance, while global explanations aim to uncover the overall functioning of a complex model [12]. Considering various purposes of explanation, the explanation techniques mainly showcase the explanation from two perspectives, including feature attributions and counterfactual examples. Feature attribution aims to provide users with important scores for each feature’s impact on model predictions, while counterfactual examples aim to offer alternative instances that explicitly assist users in grasping the model’s decision-making process. In recent years, with the growing proficiency and wide usage of black-box LLMs, especially closed-source LLMs service, post-hoc explanations have become increasingly prominent and have garnered significant attention in NLP research due to the inaccessibility of LLMs’ model weights and structure [46].
B.2 Explainability of LLMs
@@
The majority of explanation efforts in LLM research have centered on delivering explanations. One group of studies calculates importance scores for specific tokens [24, 17], another line of progress generates NL explanations by leveraging the pre-trained LLMs with internal model knowledge sources [21, 9, 28], the other group of work leverages LLMs themselves to generate chain-of-thought (CoT) reasoning [22, 34, 8, 8] as the self-explanations through the one feed-forward inference process. Furthermore, some studies aim to yield counterfactual explanations by pre-trained LLMs to assist users in better understanding the decision-making process from LLMs [7, 8]. Although NL explanations offer fantastic human-understandable insights than token-wise explanations, the explanations can lose their fidelity via one feed-forward inference process of pre-trained LLMs. Unreliability and non-fidelity of NL explanations are still a concern [45, 40]. Given our primary aim of producing faithful explanations, our efforts are concentrated on generating NL explanations to improve the likelihood of accurately representing the decision-making process of LLMs.
B.3 LLMs as Optimizers
@@
LLMs as optimizers is a novel paradigm, describing optimization problems in natural language and utilizing the reasoning capabilities of LLMs for optimizing [44]. Depicting optimization problems in natural language enables the optimization of diverse tasks without defining formal specifications, such as prompt optimization [44, 10, 16], agent learning [36], and model labeling [38]. Based on this optimization paradigm, our work introduces a generative explanation framework with a novel estimation method of sentence-level fidelity.
Appendix C Hyper-parameter Settings of FaithLM
@@
The hyper-parameters of FaithLM are given in Table 2. The configuration for explainers is consistent across Claude-2 and GPT-3.5-Turbo, provided that the parameters are adjustable. Likewise, the settings for predictors are uniform, with Phi-2 and Viucua-13B receiving identical hyperparameter configurations during the experiments conducted in this study.
@@
Fidelity-enhanced Optimization Optimization Steps 20 20 20
@@
Trigger-oriented Optimization Optimization Steps 50 100 100
@@
Appendix D Computation Infrastructure and Costs
D.1 Computation Infrastructure
@@
D.2 Computation Costs
The computational costs associated with xLLM primarily differ from the inference costs of local LLMs and the expenses related to API-accessed LLMs. The computational costs depend on the parameter scale and variants of LLMs used in the xLLM framework, shown in Table 4 and 5.
@@ D.2 Computation Costs
Appendix E Additional Experimental results of FaithLM
E.1 Optimization Procedure of derived explanations
@@
E.2 Robustness Analytics of Configurations in FaithLM
Figure 9: Robustness Analytics of FaithLM on different LLM settings, which are Temperature (left) and Top-p strategies (right).
E.3 Additional Optimization Curve of Explanation Trigger Prompt
@@
E.4 Additional Experiments on Diverse Domains of Dataset
We further have conducted additional experiments on one new MedMCQA dataset [32] in the healthcare domain. We evaluate the FaithLM using a fidelity assessment under Natural Language Explanation Generation settings. All experimental configurations follow the settings in Section 4.2. The experimental results are shown in the table below. We observe that FaithLM outperforms the baseline method, which is consistent with the experimental results across other domain datasets that were evaluated in our work.
@@ E.4 Additional Experiments on Diverse Domains of Dataset
E.5 Additional Experiments on Truthfulness Evaluation
We further lunch additional experiments on all evaluators with different evaluation settings to measure the relevance between derived explanations and ground-truth explanations. The evaluators assesses their relevance using a GPT-Score in GPT evaluator on a scale of one to five. If the two provided explanations are classified under the “similar content" category or receive a GPT-Score close to five, this indicates that the derived explanations are highly similar to ground-truth explanations.
@@ E.5 Additional Experiments on Truthfulness Evaluation
Appendix F Details of Evaluation Prompt Usage
We provide a listing of the evaluation prompts in Table 7 utilized in assessing the performance of FaithLM. The first row reveals the evaluation prompt on comparing the derived explanation with the ground-truth (GT) explanation in the ECQA dataset in Section 4.3; and the second row demonstrates the evaluation prompt on activating the GPT classifier and the GPT scorer for assessing contrary explanations in Section 4.5.
@@ Appendix F Details of Evaluation Prompt Usage
GPT classifier for Contrary Explanations Given a user instruction and two AI assistant responses, your job is to classify whether two responses in S1 and S2 belong to G-1, G-2, or G-3. The meaning of class is as follows: (G-1) same semantic meaning, (G-2) opposite semantic meaning, and (G-3) no relation. Judge responses holistically, paying special attention to whether two responses have the same semantic meaning. Judge responses with only ONE class label as your final answer. S1:{derived explanation}. S2:{Contrary Explanations}. Please ONLY respond in either G-1, G-2, or G-3; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
GPT scorer of Contrary Explanations Given a user instruction and two AI assistant responses, your job is to rate from ONE to FIVE to judge whether two responses in S1 and S2 have the same semantic meaning or not. A FIVE score refers to being totally the same, and ONE score refers to being totally the opposite. Judge responses holistically, paying special attention to whether two responses have the same semantic meaning. The judge responds with the rates between ONE and FIVE. S1:{derived explanation}. S2:{Contrary Explanations}. Please ONLY respond to the rate value; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
@@
Appendix G Details of Prompts Usage in FaithLM
We provide a listing of the prompts in Table 7 utilized in FaithLM in different tasks. The first row demonstrates the initial explanation trigger prompt leveraging in both fidelity-enhanced optimization and trigger-oriented optimization. The second row shows the prompt for the LLM agent to generate the contrary explanations.

### v4

@@ 4.5 Ablation Studies on contrary hint (RQ3)
The quality of contrary hints $\neg\mathcal{E}_{\text{NL}}$ determines the efficacy of FaithLM. We leverage the powerful LLMs as the LLM agent to generate contrary hints, requesting the delivery of high-quality opposite-meaning outputs from their original NL explanations. In this section, we evaluate the quality of contrary hints, aiming to observe the semantic differences between the original NL explanations and their contrary hints. To examine the quality, we employ one GPT-4o classifier and two well-trained NLI classifiers, Roberta-Large and XLNet-Large Nie et al. (2020). We leverage each classifier to distinguish whether the relationship between the “original NL explanations" and “contrary hints" belong to the category of “similar meaning (entailment)," “dissimilar meaning (contradiction)," or “non-relevant (neutral)." We follow the evaluation settings from Liu et al. (2023) on GPT-classifier with evaluation prompt provided in Table 7.
@@
The results are shown in Figure 7 and Figure 11 under the randomly sampled 100 instances from the ECQA and COPA datasets. We observe that the two NLI classifiers achieve up to 86% and 82% in the “dissimilar meanings"" category on the ECQA and COPA datasets, respectively. The results of the GPT-classifier demonstrate that the derived explanations from SelfExp obtain more non-faithful information than FaithLM, risking the LLM agent of Fidelity Evaluator in generating non-relevant information as the contrary hints. Case studies are provided in Appendix K to show the informativeness and readability of contrary hints.
@@ 5 Conclusion
In this paper, we introduce FaithLM to explain the decision-making process of LLMs, instead of providing reasoning or self-refinement feedback as model explanation. Specifically, FaithLM employs a fidelity enhancement strategy to progressively refine the fidelity of derived explanations and explanation trigger prompts. FaithLM conducts an iterative process to improve the fidelity of derived explanations. Theorem 1 establishes the contrary-hint score as a valid measure of faithfulness. Experimental results demonstrate the effectiveness of FaithLM, and better alignment with the ground-truth explanations. This suggest that the decision-making process are truly reflected. For future work, we plan to extend FaithLM in healthcare, where the needs for transparency is critical given the growing reliance on black-box LLMs.
6 Limitations
One significant limitation of FaithLM associated with the carbon emissions during the experiments. To generate better fidelity and truthfulness of the derived explanations, FaithLM requires iterating a few rounds of optimization during the explanation derivation. This leads to extra computational resources to proceed. The extra computational power is thus required for optimizing FaithLM leads to considerable energy consumption, which, in turn, results in a significant carbon footprint. As the demand for more sophisticated LLMs continues to grow, so does their environmental impact. This limitation underscores the urgent need to explore and adopt more sustainable practices and technologies in the development and learning FaithLM with fewer optimization steps to mitigate their ecological footprint.
@@ 6 Limitations
Appendix A Proof of Theorem
Theorem 1 (Latent-Context Intervention Validity for Faithfulness).
Let $f:\mathcal{X}\!\times\!\mathcal{C}\!\to\!\Delta(\mathcal{Y})$ be a language model mapping an input $X$ and latent context $C$ to a predictive distribution over an output space $\mathcal{Y}$. Let $E_{NL}$ denote a natural-language explanation of $f(X;C)$, and let $\lnot E_{NL}$ denote its contrary hint. Assume that $E_{NL}$ asserts a proposition about a semantic factor $S_{E}=s(X,C)$, where $s(\cdot)$ extracts the decision-relevant concept, which is latent or retrieved, that the explanation verbalizes. Conditioning on $\lnot E_{NL}$ is equivalent to intervening on this factor while holding $(X,C)$ fixed, i.e., $f(X;C\mid\lnot E_{NL})=f(X;C\mid do(S_{E}\!\leftarrow\!\bar{s}))$ for some contradictory value $\bar{s}$, and predictions are invariant to any irrelevant text $R$, so $f(X;C)=f(X\cup R;C)$. Defining $S_{E}(X;C)=D(f(X;C),f(X;C\mid\lnot E_{NL}))$, where $D$ is any strictly proper divergence, we have
$S_{E}(X;C)=0$ $\iff E_{NL}\text{ is non-faithful for }f(X;C),$
$S_{E}(X;C)>0$ $\iff E_{NL}\text{ is faithful for }f(X;C).$
Hence, the contrary-hint score $S_{E}$ constitutes a valid empirical estimator of causal faithfulness even when the decision-relevant content is not contained in the observed input $X$ but arises from latent or retrieved context.
Proof.
By the causal definition of faithfulness, an explanation is faithful if and only if altering the truth value of the decision-relevant semantic factor changes the model’s prediction while holding other causes fixed. Because $E_{NL}$ asserts a proposition about $S_{E}=s(X,C)$, and conditioning on $\lnot E_{NL}$ implements $do(S_{E}\!\leftarrow\!\bar{s})$ with $(X,C)$ fixed, we have $f(X;C\mid\lnot E_{NL})=f\!\big(X;C\,\big|\,do(S_{E}\!\leftarrow\!\bar{s})\big)$. If $E_{NL}$ is non-faithful, then $S_{E}$ has no causal influence on $Y$ under $(X,C)$ and the intervention leaves the predictive distribution unchanged, yielding $S_{E}(X;C)=0$. If $E_{NL}$ is faithful, $S_{E}$ lies on a causal path to $Y$ under $(X,C)$ and the intervention changes the predictive distribution; strict propriety of $D$ implies $S_{E}(X;C)>0$. Invariance under irrelevant spans follows by the stated stability condition and applies to both terms inside $D$. ∎
Corollary 1 (Robustness and Monotonicity).
Under contextual stability, the contrary-hint score remains invariant to irrelevant input variations, satisfying
$S_{E}(X;C)=S_{E}(X\cup R;C)\quad\text{for any semantically irrelevant }R.$
Furthermore, if an iterative procedure produces a sequence of explanations $\{E_{NL}^{(t)}\}$ such that $S_{E}^{(t+1)}(X;C)\geq S_{E}^{(t)}(X;C)$ for all iterations prior to convergence or a decision flip, then the sequence is non-decreasing in causal faithfulness. Consequently, any iteration with $S_{E}^{(t)}(X;C)>0$ corresponds to a faithful explanation.
Appendix B Details about Datasets
@@
  • ECQA Aggarwal et al. (2021). ECQA is an extension of the CQA dataset Talmor et al. (2019). Specifically, based on the CQA dataset, it annotates the positive or negative properties and golden explanations for the QA pairs. Due to API cost budgets, we evaluate our framework on the first 500 instances in the ECQA dataset.
@@
  • TriviaQA LongBench Joshi et al. (2017). TriviaQA LongBench (TriviaQA-Long) is a reading comprehension dataset. It includes 300 question-answer-evidence triples sourced from the Longbench Bai et al. (2023) dataset [^2] https://huggingface.co/datasets/THUDM/LongBench/. This dataset features question-answer pairs crafted by trivia enthusiasts, accompanied by independently sourced evidence documents, providing supervision for answering these questions.
@@
  • Balanced COPA Roemmele et al. (2011); Kavumba et al. (2019). The Balanced COPA (COPA) dataset is a collection of 500 questions for commonsense causal reasoning. Each question consists of a premise and two alternatives, where the task is to select the alternative that more plausibly has a causal relation with the premise.
Appendix C Experiment Settings
We introduce the experimental settings for evaluating FaithLM. Two distinct types of explanation tasks and evaluation settings are as follows.
Fidelity-enhanced Explanation
In this task, our goal is to produce NL explanations that exhibit a higher fidelity. The fidelity is exploited as a metric to evaluate fidelity. FaithLM is evaluated across all testing instances, where an NL explanation is generated for each instance, and the averaged fidelity score is calculated, serving as the reported metric to evaluate fidelity.
Explanation Trigger Prompt Optimization.
Evaluation Metrics. The quality of the derived NL explanation is evaluated under the fidelity and truthfulness metrics. The fidelity follows the prior work Chen et al. (2025), which observes the discrepancy of the targeted LLMs by incorporating contrary hints to the input. The evaluation of truthfulness assesses the correlation between the derived NL explanations to the ground-truth explanations Specifically, we leverage GPT-4o and two well-trained natural language inference (NLI) models, Roberta-Large and XLNet-Large Nie et al. (2020) from the huggingface hub Wolf et al. (2019), as the evaluators. With the same evaluators setup, the truthfulness evaluation follows the settings from Liu et al. (2023), and uses the evaluation prompt provided in Appendix H. Specifically, the evaluators assess the derived explanations and ground-truth explanations, determining whether the two sentences belong to “similar content", “dissimilar content," or “non-relevant content". Higher the proportion of “similar content", the more consistent results with ground-truth NL explanations.
Implementation Details. In the experiments, we explore two variants of LLMs as the targeted LLMs $f(\cdot)$: Vicuna-7B Chiang et al. (2023) and Phi-2 Javaheripi and Bubeck (2023), two types of LLMs as the explainers $\textsl{g}(\cdot)$ in FaithLM: GPT-3.5-Turbo and Claude-2 Anthropic (2023). The LLM agent for generating the contrary hints takes the same LLMs as those used by the explainers. All reported results are calculated from the average scores of 3 times repetitions with the grid search on the performance. The settings for predictors are uniform, with Phi-2 (2.7B) and Viucua-7B receiving identical hyperparameter configurations during the experiments conducted in this study.
Appendix D Related Work
D.1 Post-hoc Explanation
@@
Post-hoc explanation techniques have undergone significant development and discussion, driven by the widespread adoption of black-box ML models across various data modalities. A multitude of post-hoc algorithms has been introduced from two aspects: local and global explanations Molnar (2022); Du et al. (2019). Explanations aim to explain the reasoning behind an individual model for each input instance, while global explanations aim to uncover the overall functioning of a complex model Chuang et al. (2023). Considering various purposes of explanation, the explanation techniques mainly showcase the explanation from two perspectives, including feature attributions and counterfactual examples. Feature attribution aims to provide users with important scores for each feature’s impact on model predictions, while counterfactual examples aim to offer alternative instances that explicitly assist users in grasping the model’s decision-making process. In recent years, with the growing proficiency and wide usage of black-box LLMs, especially closed-source LLMs service, post-hoc explanations have become increasingly prominent and have garnered significant attention in NLP research due to the inaccessibility of LLMs’ model weights and structure Zhao et al. (2023).
D.2 Explainability of LLMs
@@
The majority of explanation efforts in LLM research have centered on delivering explanations. One group of studies calculates importance scores for specific tokens Lopardo et al. (2023); Huang et al. (2023), another line of progress generates NL explanations by leveraging the pre-trained LLMs with internal model knowledge sources Kumar and Talukdar (2020); Chen et al. (2023c); Menon et al. (2023), the other group of work leverages LLMs themselves to generate chain-of-thought (CoT) reasoning Lanham et al. (2023); Radhakrishnan et al. (2023); Chen et al. (2023b, b) as the self-explanations through the one feed-forward inference process. Furthermore, some studies aim to yield counterfactual explanations by pre-trained LLMs to assist users in better understanding the decision-making process from LLMs Chen et al. (2021, 2023b). Although NL explanations offer fantastic human-understandable insights than token-wise explanations, the explanations can lose their fidelity via one feed-forward inference process of pre-trained LLMs. Unreliability and non-fidelity of NL explanations are still a concern Ye and Durrett (2022); Turpin et al. (2023). Given our primary aim of producing faithful explanations, our efforts are to generate NL explanations to improve the likelihood of accurately representing the decision-making process of LLMs.
D.3 LLMs as Optimizers
@@
LLMs as optimizers is a novel paradigm, describing optimization problems in natural language and utilizing the reasoning capabilities of LLMs for optimizing Yang et al. (2023). Depicting optimization problems in natural language enables the optimization of diverse tasks without defining formal specifications, such as prompt optimization Yang et al. (2023); Cheng et al. (2023); Guo et al. (2023), agent learning Shinn et al. (2023), and model labeling Thomas et al. (2023). Based on this optimization paradigm, our work introduces a generative explanation framework with a novel estimation method of sentence-level fidelity.
Appendix E Hyper-parameter Settings of FaithLM
@@
The hyper-parameters of FaithLM are given in Table 2. The configuration for explainers is consistent across Claude-2 and GPT-3.5-Turbo, provided that the parameters are adjustable. Likewise, the settings for predictors are uniform, with Phi-2 and Viucua-7B receiving identical hyperparameter configurations during the experiments conducted in this study.
@@
Fidelity-enhanced
Optimization
Optimization Steps 20 20 20
@@
Trigger-oriented
Optimization
Optimization Steps 50 100 100
@@
Appendix F Computation Infrastructure and Costs
F.1 Computation Infrastructure
@@
F.2 Computation Costs
The computational costs associated with FaithLM primarily differ from the inference costs of local LLMs and the expenses related to API-accessed LLMs. The computational costs depend on the parameter scale and variants of LLMs used in the FaithLM framework, shown in Table 4 and 5.
@@ F.2 Computation Costs
Appendix G Additional Experimental results of FaithLM
G.1 Optimization Procedure of derived explanations
@@
G.2 Additional Optimization Curve of Explanation Trigger Prompt
@@
G.3 Additional Experiments on Diverse Domains of Dataset
We further have conducted additional experiments on one new MedMCQA dataset Pal et al. (2022) in the healthcare domain. We evaluate the FaithLM using a fidelity assessment under Natural Language Explanation Generation settings. All experimental configurations follow the settings in Section 4.2. The experimental results are shown in the table below. We observe that FaithLM outperforms the baseline method, which is consistent with the experimental results across other domain datasets that were evaluated in our work.
@@ G.3 Additional Experiments on Diverse Domains of Dataset
G.4 Additional Experiments on Truthfulness Evaluation
We evaluate all baseline methods and FaithLM under multiple settings to assess how closely the derived explanations match the ground-truth rationales. A GPT-based evaluator assigns a GPT-Score from 1 to 5, with higher values indicating greater semantic similarity. Explanations tagged as “similar content” or scoring near 5 are treated as matches. We report the truthfulness of generated explanations in Figure 10. A GPT-based evaluator assigns a truthfulness score from 1 to 5 by checking factual consistency with the input, the task label, and commonsense knowledge. Explanations tagged as factually consistent or scoring near 5 are counted as truthful. Across all settings, FaithLM attains the highest mean truthfulness score and the largest fraction of truthful explanations, indicating that our optimization improves not only fidelity but also factual quality of the produced rationales.
@@ G.4 Additional Experiments on Truthfulness Evaluation
G.5 Additional Experiments of Contrary Hints.
We here showcase the results of using NLI classifiers to evaluate the quality of contrary hints.
Figure 11: Quality Evaluation results of Contrary Hints via NLI classifiers
Appendix H Details of Evaluation Prompt Usage
We provide a listing of the evaluation prompts in Table 7 utilized in assessing the performance of FaithLM. The first row reveals the evaluation prompt on comparing the derived explanation with the ground-truth (GT) explanation in the ECQA dataset in Section 4.3; and the second row demonstrates the evaluation prompt on activating the GPT classifier and the GPT scorer for assessing contrary hints in Section 4.5.
@@ Appendix H Details of Evaluation Prompt Usage
GPT classifier for contrary hints Given a user instruction and two AI assistant responses, your job is to classify whether two responses in S1 and S2 belong to G-1, G-2, or G-3. The meaning of class is as follows: (G-1) same semantic meaning, (G-2) opposite semantic meaning, and (G-3) no relation. Judge responses holistically, paying special attention to whether two responses have the same semantic meaning. Judge responses with only ONE class label as your final answer. S1:{derived explanation}. S2:{contrary hints}. Please ONLY respond in either G-1, G-2, or G-3; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
GPT scorer of contrary hints Given a user instruction and two AI assistant responses, your job is to rate from ONE to FIVE to judge whether two responses in S1 and S2 have the same semantic meaning or not. A FIVE score refers to being totally the same, and ONE score refers to being totally the opposite. Judge responses holistically, paying special attention to whether two responses have the same semantic meaning. The judge responds with the rates between ONE and FIVE. S1:{derived explanation}. S2:{contrary hints}. Please ONLY respond to the rate value; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.
@@
H.1 Robustness Analytics of Configuration (RQ3)
Figure 12: Robustness Analytics of FaithLM: Temperature (left) and Top-p strategies (right).
Appendix I Details of Prompts Usage in FaithLM
We provide a listing of the prompts in Table 7 utilized in FaithLM in different tasks. The first row demonstrates the initial explanation trigger prompt leveraging in both fidelity-enhanced optimization and trigger-oriented optimization. The second row shows the prompt for the LLM agent to generate the contrary hints.